# How AI Ideas Affect the Creativity, Diversity, and Evolution of Human Ideas: Evidence From a Large, Dynamic Experiment

### v2

@@ 6.2. Global
We used a Monte Carlo procedure and permutation tests to assess if conditions differed with respect to these metrics. For 50 Monte Carlo runs, for each {[condition], item} combination, we randomly sampled 50 ideas and computed idea diversity metrics. We then conducted pairwise paired (at the level of Monte Carlo seeds and items) permutation tests with 10,000 iterations to see if the two conditions differed on these metrics. As a non-parametric measure of effect size, we also calculate Cliff’s Delta $(\delta)$, which ranges from -1 to 1. A value of 0 indicates no difference between the two conditions, +1 indicates values from the first condition are always larger, and -1 indicates the opposite. See Appendix I.1 for more details.
6.3. Evolution
6.3.1. Creativity
@@
6.3.2. Idea Diversity
Intuitively, we are interested in if—as the experiment goes on—ideas that participants submit tend to become more or less similar to each other. We use the trial number in a response chain to index time in the experiment. For example, is the set of submitted responses at trial number 4 more or less similar to each other as the set of submitted responses at trial number 20? Here, the diversity of interest is not between a submitted response and example responses but between all submitted responses at a given ‘time point’ (i.e., trial number). The question is if the diversity increases or decreases as the experiment goes on and if this rate of change differs by condition. Here is the mechanics of our process. See Appendix I.3 for more details.
@@ 6.3.2. Idea Diversity
7. Results
7.1. Creativity
We found no effect of conditions on creativity. Average individual creativity did not significantly differ by condition $(F(4,19.86)=0.12,p=0.97)$ and no condition coefficient differed from zero in our regression (Appendix Table 19). Hence, we conclude that neither AI exposure nor AI disclosure affected individual creativity. Additionally, we tested for whether the evolution of creativity differed by condition via a likelihood ratio test on whether interacting trial number and experimental condition would improve the model fit. The likelihood ratio test indicated that allowing for these interactions did not significantly improve the model fit $(\chi^{2}(4)=6.52,p=0.16)$ [^11] However, there was a small, negative interaction effect ($\beta=-0.015$, $t(3248)=-2.21$, 95% CI = $[-0.03,-0.002]$) between trial number and the [High Exposure, Disclosed] condition when adding this interaction. However, due to the (1) size of the interaction combined with (2) no overall differences and (3) a null likelihood ratio test, we do not interpret this interaction.. In short, we do not find enough evidence to conclude creativity was affected by experimental conditions.
7.2. Idea Diversity
7.2.1. Local Level
Intuitively, local idea diversity is how different a response is from the examples a participant sees. There was no main effect of condition $(F(4,19.95)=0.09,p=0.98)$, and the effect of self-perceived creativity did not differ by condition $(F(4,2650.44)=1.59,p=0.18)$. But the effect of belief in AI’s relative creativity did differ by condition, $F(4,2635.80)=2.93,p=0.02$. As robustness checks, we ran the same specification with two alternative measures of idea diversity, mean pairwise distance and distance from the centroid. Regression results are broadly similar. However, post-hoc estimated marginal means showed non-significant contrasts, so we refrain from interpreting this finding. See Appendix I.2 for a more in-depth discussion, regression results, and pairwise comparisons.
7.2.2. Global
Figure 5. Median pairwise distance of submitted ideas in a condition. There was more global diversity of ideas in the high AI exposure conditions than in the control condition.
[…]
7.2.3. Evolution
Figure 6. High exposure to AI ideas increased the rate of change in idea diversity.
By measuring the evolution of idea diversity, we capture the rate of change in idea diversity across trials. See Appendix I.3 for more details on the procedure. Relative to the control condition, the conditions with high exposure to AI ideas (but not low exposure to AI) had increased rates of change in idea diversity. See Figure 6 for estimated marginal means predictions and Appendix Table 18 for regression results. As with global idea diversity, different metrics yielded similar regression coefficients. In the control condition, idea diversity decreased over trials ($\beta=-0.39$, $t(349)=-2.23$, 95% CI = $[-0.73,-0.05]$, $p=0.03$). That is, submitted ideas were becoming more similar to each other as the experiment went on. Relative to the control condition, however, the slope of idea diversity with respect to trial number was more positive for the [High Exposure, Undisclosed] condition ($\beta=0.53$, $t(349)=2.2$, 95% CI = $[0.06,0.99]$, $p=0.03$) and the[High Exposure, Disclosed] condition ($\beta=0.57$, $t(349)=2.37$, 95% CI = $[0.1,1.03]$, $p=0.02$). The rate of change in idea diversity for the low AI exposure conditions did not differ from the rate of change in the control condition. Thus, we conclude that high exposure to AI ideas increased the rate of idea diversity relative to the no-AI, control condition.
7.3. AI Adoption
7.3.1. Local Level
At the local level, we measured AI adoption by the maximum cosine similarity between a participant’s response and AI examples the participant saw. There was a main effect of condition $(F(3,16.59)=4.33,p=0.02)$. But we would expect higher similarity to AI ideas in the high-exposure condition even by chance (since there are more AI ideas), so we do not interpret main effects and instead focus on subgroup differences and effects of disclosure in the high-exposure condition. We found that the effect of conditions did not differ by interest groups $(F(6,719.77)=1.98,p=0.07)$, but the effect of conditions did differ by self-perceived creativity $(F(3,1984.95)=5.18,p=0.001)$ and relative AI creativity $(F(3,1974.58)=2.9,p=0.03)$. As robustness checks, we ran the same specification with two alternative measures of AI adoption, mean and median AI adoption. The coefficients of our regression are broadly similar. See Appendix K for regression results and post-hoc contrasts.
Figure 7. In the [High Exposure, Disclosed] condition, participants high in self-reported creativity had higher AI adoption (raw data).
Figure 8. Adoption of AI ideas by self-perceived creativity In the [High Exposure] conditions (estimated marginal means).
Figure 9. Higher creativity participants adopted ideas solely based on exposure, not disclosure (estimated marginal means).
Exposure to AI ideas increased adoption for (self-perceived) high-creativity participants regardless of disclosure, but this was not the case for (self-perceived) low-creativity participants. There was a significant interaction between self-perceived human creativity and the [High Exposure, Disclosed] condition ($\beta=0.11$, $t(2588)=3.93$, 95% CI = $[0.06,0.17],p=0.0001$; Appendix Table 23). To probe this interaction, we used our model to predict AI adoption by condition for both the top 10% and bottom 10% of participants by self-perceived creativity (Figures 9, 9 and 9). For high-creativity participants (Figure 9), adoption rates appear to differ only by exposure (color) and not disclosure (shape). More formally, we tested whether the effect of exposure on adoption is larger when AI ideas are disclosed vs undisclosed. We find that for high-creativity participants, there is no difference in adoption between ([High Exposure, Undisclosed] - [Low Exposure, Undisclosed]) and ([High Exposure, Disclosed] - [Low Exposure, Disclosed]), $\Delta=-1.69,d=-0.14,p=0.59$. That is, the effect of exposure is not moderated by disclosure. But for low-creativity participants, the difference in adoption for the undisclosed conditions ([High Exposure, Undisclosed] - [Low Exposure, Undisclosed]) was larger than the equivalent difference in adoption for disclosed conditions ([High Exposure, Disclosed] - [Low Exposure, Disclosed]), $\Delta=7.77,d=0.65,p=0.03$. That is, disclosing ideas as from AI reduced the effect of exposure on adoption for lower (self-reported) creativity participants. In summary, higher (self-reported) creativity people adopt AI ideas solely based on content, and not disclosure.
We also found that one’s attitude about AI’s creativity affected adoption, though this had a smaller effect than self-perceived creativity. There was a significant interaction between the [High Exposure, Undisclosed] condition and relative AI creativity (positive values imply AI is more creative than humans), $\beta=-0.07$, $t(2588)=-2.61$, 95% CI = $[-0.13,-0.02],p=0.01$. We used estimated marginal means to probe this interaction by predicting AI adoption for the top and bottom decile of participants by belief in relative AI creativity. We found that in the [High Exposure, Undisclosed] condition, people who believed AI was uncreative (bottom decile of AiRelCreate) were slightly more likely to adopt AI ideas than people who believed AI was creative (top decile of AiRelCreate), $\Delta=4.66,d=0.39,p=0.005$. But no such difference existed in the [High Exposure, Disclosed] condition. This may suggest labeling sources as AI neutralizes adoption among users who do not think AI is creative.
Figure 10. Adoption of AI ideas in the [High Exposure, Disclosed] condition versus ‘difficulty’ of prompt. When ideas were disclosed as from AI, participants adopted AI ideas for difficult prompts.
In addition to who adopts AI ideas, we also measured when AI ideas are adopted. We found that people adopt AI ideas for difficult prompts rather than easier prompts (Figure 10). To measure the ‘difficulty’ of an item prompt, we first calculated the mean creativity of a response to an item in the control condition. Then we reverse-ranked items such that high mean creativity implies low difficulty and vice versa. We measured ‘AI adoption‘ of an item by the average of the trial-level maximum similarity to AI examples. We then examined the rank-order correlation between item difficulty and AI adoption in high-exposure conditions. If task difficulty leads people to rely on AI, then we should see a larger correlation between item difficulty and AI adoption in the [High Exposure, Disclosed] condition than in the [High Exposure, Undisclosed] condition. That is what we find. The rank-order correlation between difficulty and adoption was $\rho=0.8$ for the [High Exposure, Disclosed] condition but only $\rho=0.3$ for the [High Exposure, Undisclosed]. That is, when people were told ideas were from AI, they were more likely to adopt AI ideas if the prompt was difficult. Since we employed only five items, we view this finding as speculative; future work should test this relationship with a larger number of stimuli.
8. Discussion
Against the backdrop of a massive increase in LLM exposure, we asked: How does exposure to ideas generated by LLMs affect the creativity, diversity, and evolution of human ideas? To answer this, we conducted a large-scale experiment where participants submitted ideas in response to the Alternate Uses Task (a measure of creativity where people brainstorm novel uses of an item) after seeing a set of example ideas. The examples were from prior participants in the same experimental condition or—in some conditions—ChatGPT. The evolving aspect of our experiment, that ideas in a condition feed forward to subsequent trials in that condition, captures the interdependent nature of idea formation and lets us model the evolutionary effects of having AI ‘in the culture loop’. Here are three takeaways from our experiment.
8.1. AI makes ideas different but not better.
On the other hand, the presence of AI ideas increased the diversity of human ideas. This is consistent with work suggesting collaborating with AI leads to more diverse or unexpected outputs (Yang et al., 2022; Osone et al., 2021; Lee et al., 2022; Branch et al., 2021; Gero and Chilton, 2019) and inconsistent with other work that finds collaborating with LLMs decreases diversity (Padmakumar and He, 2024; Doshi and Hauser, 2024; Dell’Acqua et al., 2023). But we highlight that our study is testing passive exposure to AI ideas and not active engagement with AI ideas. Our setup—passive exposure to AI ideas, scattered amongst human ones—maps onto how many users now experience AI ideas. Hence, it may be that active engagement with LLMs decreases content diversity but simply seeing these ideas as ‘sparks’ (Gero et al., 2022) increases content diversity. And because many more people may be passively exposed to LLM outputs than actively engaging with LLMs, the effect of passive exposure is important to understand.
Crucially, high AI exposure increased both average amounts of diversity and rates of change in idea diversity. The latter result is especially important. Small differences in rates of change can yield large aggregate differences over time. Future work—both simulations and dynamic experiments—can explore the implications of this increase in collective idea diversity unaccompanied by an increase in average individual creativity. For instance, can this dynamic generate ‘innovation’?
Our finding around the evolution of diversity (Figure 6) is instructive. Seeing other people’s ideas reduced idea diversity in the control condition over time. This may suggest that successive participants were converging on a particular idea sequence. But then injecting AI ideas into the example set increased the diversity of submitted responses by partially ‘resetting’ this convergence. Our finding relates to recent work proposing AI systems that generate ‘alien’ scientific hypotheses humans would not think of (Sourati and Evans, 2023). More generally, a promising avenue for future work: Can AI input reduce ‘groupthink’?
8.2. High creativity people are less influenced by the source label of ideas.
Participants who viewed themselves as highly creative had the same levels of adoption of AI ideas in both disclosed and non-disclosed conditions. But for lower-creativity participants, knowing the source of an idea did affect the adoption of that idea. Perhaps people high in self-reported creativity relied less on source cues when adopting ideas because they were more confident in their ability to judge an idea’s creative merit. Future work can employ think-alouds to better understand how AI disclosure affects the idea-generation process, itself. Regardless, our results suggest that (self-reported) creative people will adopt ideas on the basis of their content. Knowing the source does not matter. In a world where humans have difficulty distinguishing if the content was human or AI-generated (Jakesch et al., 2022), these findings suggest people high in (self-reported) creativity will not be ‘duped’ into adopting AI ideas.
8.3. Participants may adopt AI ideas when the task is difficult
When AI ideas were labeled, participants were more likely to adopt AI ideas for difficult prompts rather than easy prompts. Although this finding should be taken as speculative (due to the small number of items), it is similar to what (Roemmele, 2021) observed, where seeing AI examples only influenced creative output when the task was difficult. Both our and (Roemmele, 2021)’s results are consistent with a theoretical account of task difficulty being associated with increased reliance on automation (Goddard et al., 2014). Future work can further test whether users adopt AI ideas for more challenging creative tasks.
If users turn to AI for difficult rather than trivial tasks, this has several implications. On one hand, AI can augment human creativity where human imaginations falter. At the same time, researchers raised concerns over ‘model collapse’ (Shumailov et al., 2023)—the deteriorating performance of LLMs when trained on their outputs. If reliance on AI for creative tasks becomes routine, this may contribute to model collapse, ironically decreasing the efficacy of such reliance.
8.4. Conclusion: Passive exposure to AI ideas affects collective thought.
We conclude that passive exposure to AI ideas—the kind of passive exposure we are inundated with in a post-ChatGPT era—does affect collective thought. Even small effects are meaningful since this exposure is both pervasive and growing. But the effects of AI ideas are nuanced. Seeing AI ideas did not increase individual creativity, though it did increase collective diversity. The effects of AI ideas vary across individuals and tasks. There is still much to learn. We hope our study inspires more research on how passive exposure to AI ideas affects collective thought.
9. Limitations & Future Work
Our study has several limitations that can inform future work. First, we measured the effect of AI ideas for a single task. We chose this task because it is one of the most common creativity tasks (Abraham, 2016). But future work could explore if our results replicate for other kinds of tasks. Second, we had to operationalize ‘ChatGPT’ in some concrete way. The logic for our prompt was driven by ecological validity and prior work: We used a zero-shot prompt because that is what users would likely use, and the specific prompt we used was derived from prior research. We chose not to vary prompts in order not to further increase the complexity of an already complex experiment. Future work could explore if different prompts elicit different results. Another avenue for future work is only propagating the ‘best’ AI ideas forward. Third, future work should test if alternative classifiers or ways of conceiving variables yield different results. For idea diversity and AI adoption, we addressed this problem by showing that conceptually similar ways of measuring variables yielded qualitatively similar results. For our creativity measure, we used a highly accurate classifier (correlation with human judgments greater than $r=0.88$ for items we used) trained for this exact task, for these exact items. But of course, all models have some error and future work based on this model propagates these errors. Incidentally, human judges of creativity only correlate with other human judges at $r=0.88$ (Organisciak et al., 2023), suggesting the classifier we used may be approaching ‘the approximate ceiling at which we could expect a model to correlate with human judgments’ (Organisciak et al., 2023, pg. 11) of creativity. Nonetheless, future work can adopt a similar design but with human judgments or different measures. Fourth, our finding about AI adoption and task difficulty is based on five AUT items. Future work should explore this relationship with a larger number of stimuli. Fifth, we focus on one facet of creativity: originality. Future work can also explore whether AI ideas have different effects on other facets of creativity. We also did not necessarily measure idea “quality”, which is distinct from originality and diversity. Sixth, we employed a convenience sample of technology-interested users and creative professionals. While these two groups are most relevant to the phenomena in question, our sample also limits generalizability. Future work can explore these dynamics with different samples. Seventh, these dynamics may differ for future LLMs. Finally, we conducted this experiment close to the launch of ChatGPT. As AI becomes increasingly embedded in everyday life, attitudes towards AI and ways of engaging with AI may also change. Despite these limitations, our work offers the first large-scale, dynamic account of how ideas from LLMs affect collective thought.
Appendix A AUT Items
@@
Table 5. AUT items by frequency of occurrence in dataset and classifier accuracy. Accuracy is defined as the correlation between human ratings of creativity and model predictions. The overall accuracy was r = 0.81. The accuracy of responses from the best-performing 5-item subset was r= 0.90. The data and model is from Organisciak et al. (2022).
@@
Appendix B AUT Prompts
We conducted a Monte Carlo experiment to confirm our modified zero-shot prompt resulted in responses with a similar word length to humans. Using parameters from Stevenson et al. (2022)’s experiments: For $n=1000$ trials, we fixed presence penalty and frequency penalty at 1, randomly chose a temperature (higher values lead to more randomness) in [0.65, 0.7, 0.75, 0.8], and randomly chose one of our 5 AUT items. For each trial, ChatGPT generated five ideas. The modified prompt resulted in responses with an average word length (M=4.44, SD = 1.34) much closer to human responses (M=4.56, SD = 4.97) than the original zero-shot prompt (M=25.38, SD=8.55). A permutation test further shows that this difference in word count was significant at $p<0.001$. We used the ideas generated by our modified zero-shot prompt as stimuli for the main experiment.
Table 6. Summary statistics of AUT prompt experiment. Human ideas are from Organisciak et al. (2022) and include only those ideas in response to the chosen AUT items. Note that in some cases ChatGPT did not return the desired number of ideas, leading to a slight discrepancy between ideas generated between the two prompts.
@@ Appendix B AUT Prompts
Appendix C Pre-Treatment Questions
Pew asked about feeling towards AI (Nadeem, 2023, 2022) and we used the specific phrasing and choice ordering from (Nadeem, 2022). We randomized the first two options and kept neutral last. Our gender question was based on guidance from Spiel et al. (2019). The options were: ’woman’, ’man’, ’non-binary’, ’prefer to self-describe’, ’prefer not to disclose’. We added a text box meant for those who preferred to self-describe. The only deviation from Spiel et al. (2019) is that we did not allow for participants to select multiple options. We note that gender (as well as age and country) were optional.
Appendix D Exclusion Criteria for Analysis
Participants could have consented and answered pre-treatment questions but failed to complete any trial. We only analyze data from participants who completed at least one trial. In $(n=4)$ cases, users submitted ages that were implausible. We replaced these age values with missing for the purpose of summarizing participants but kept the responses. In $(n=2)$ cases, responses that should not have been shown were shown. We remove these two responses from analysis. As discussed in L, we instituted content moderation after receiving several troll responses. After the study, we manually inspected each response flagged by our system. There were $46$ ideas labeled as profane, and we determined $36$ were true positives. We remove the true positives $(n=36)$ from analysis, resulting in a final set of $3414$ responses for analysis from an initial set of $3452$ responses. Importantly, we conducted chi-squared tests and found that condition was unrelated to the number of flagged ideas ($\chi^{2}$(4) = 6.06, p = 0.19), number of flagged ideas minus false positives ($\chi^{2}$(4) = 2.92, p = 0.57) or total number of excluded ideas ($\chi^{2}$(4) = 3.87, p = 0.42).
Appendix E Human vs AI Ideas
We compared a sample of 1500 ideas from our modified Stevenson prompt in the prompt experiment and a random sample of 1500 ideas from the Organisciak Dataset for our 5 items. For each set, we used the model’s predicted originality scores. Originality ranges from 1-5. Overall, ChatGPT ideas had higher ($\beta=0.62$, $t(2994)=22.49$, 95% CI = $[0.56,0.67]$) originality.
Table 7. Comparing predicted originality of ChatGPT generated ideas to ideas from a dataset of prior human responses
@@ Appendix F Participant Feedback
Additionally, we wanted to minimize attrition for participants once they started. We gave participants two pieces of feedback after each trial so they would continue taking the study. See M for screenshots.
@@ Appendix H Model Selection
Appendix I Idea Diversity
I.1. Global
For 50 Monte Carlo runs with different seed values, we sampled 50 ideas for each {[condition], item} combination. For each 50-idea set, we computed various idea diversity measures. First, we calculated all pairwise SBERT distances. Next, we measured the mean and median pairwise distances. We also computed the centroid of each 50-idea set and calculated the mean distance from the centroid. After calculating these metrics, we conducted two-tailed, paired permutation tests (10,000 iterations) to test if two conditions differed on these metrics. To conduct the paired permutation test, we randomly swapped the sign of the difference between pairs of values, equivalent to randomly swapping the condition labels of rows—simulating the null hypothesis that conditions do not differ. We then counted the proportion of null distribution iterations where one would observe a larger absolute difference in means than the observed difference. We added a 1 to the numerator and denominator, which is a common, conservative adjustment (Ojala and Garriga, 2009) and stops p-values from being 0. Because the test is paired (equivalent to swapping the condition label within each ‘row’), our permutation tests are controlling for both AUT items and Monte Carlo seeds, since each row shares these attributes. We controlled for multiple pairwise comparisons by applying a Holm-Bonferroni adjustment to p-values. As a non-parametric measure of effect size, we used Cliff’s Delta. This metric ranges from -1 to +1 where 0 indicates no difference between conditions, +1 indicates that all Monte Carlo runs for the first condition are larger than those for the second, and vice versa for -1. As with evolution, to avoid the confounding effect of conditions differing in the number of seeds, we consider ideas after the sixth trial (see SI I.3 for a more detailed discussion). This is because the experiment is designed to ‘shed’ all seed examples after trial six.
@@ I.1. Global
I.2. Local
We found mixed evidence that belief in AI’s relative creativity moderates local idea diversity. Regression results showed a small but significant interaction effect between the [High Exposure, Undisclosed] condition and relative AI creativity ($\beta=0.038$, $t(3244)=2.149$, 95% CI = $[0.003,0.073]$, $p=0.03$). We probed this effect with estimated marginal means, predicting local idea diversity for the bottom and top decile of participants by perception of AI creativity. Top-decile participants had slightly higher local idea diversity than bottom-decile participants in the [High Exposure, Undisclosed] condition ($\Delta=2.62,d=0.34$) but although this difference was significant before multiple comparisons ($p=0.01$), it was not significant after adjusting for multiple comparisons, ($p=0.06$; see Appendix Table 17). Hence, we conclude there is mixed evidence for the role of belief in AI’s relative creativity as a moderator of local idea diversity.
@@ I.3. Evolution
We took two additional steps to make sure our results were not driven by confounding factors. First, $Nobs$ controls for how many ideas are in the set that is being analyzed. Recall that we designed the experiment so that each (item, condition) combination was replicated exactly seven times in response chains of exactly 20 trials. However, there were some minor deviations (discussed in SI L) in response chains, resulting in some (item, condition, trial number) sets having fewer items than others. Hence, we control for the number of ideas in a set. Second, we only ran this analysis on data after the sixth trial in a response chain to rule out the effect of seeds on evolution. The logic here is that the condition with the most initial seeds (6) was the control condition. The experiment is designed to ‘shed’ all seeds after trial six since by that time there would have been six experiment responses, meaning the most recent six ideas in the control condition would now all be from the experiment, and hence no seeds present in the example sets. (Note that for all local analyses, we directly control for the number of seeds present in the example set as a fixed effect.)
@@
Appendix K AI Adoption
@@
Table 21. Estimated marginal means contrasts of AI adoption, using a mixed model to compare predictions for the top 10 percentile and bottom 10 percentile of participants by belief in relative AI creativity. This metric captures how creative participants think AI is relative to humans (higher values means more creative than humans). AI Adoption is the max cosine similarity of a participant’s response and AI examples. P-values adjusted for multiple comparisons using Holm-Bonferroni method.

### v3

@@ B.2. Global
We used a Monte Carlo procedure and permutation tests to assess if conditions differed with respect to these metrics. For 50 Monte Carlo runs, for each {[condition], item} combination, we randomly sampled 50 ideas and computed idea diversity metrics. We then conducted pairwise paired (at the level of Monte Carlo seeds and items) permutation tests with 10,000 iterations to see if the two conditions differed on these metrics. As a non-parametric measure of effect size, we also calculate Cliff’s Delta $(\delta)$, which ranges from -1 to 1. A value of 0 indicates no difference between the two conditions, +1 indicates values from the first condition are always larger, and -1 indicates the opposite. See SI 8.1 for more details.
B.3. Evolution
Creativity
@@ B.3. Evolution
Idea Diversity
Intuitively, we are interested in if—as the experiment goes on—ideas that participants submit tend to become more or less similar to each other. We use the trial number in a response chain to index time in the experiment. For example, is the set of submitted responses at trial number 4 more or less similar to each other as the set of submitted responses at trial number 20? Here, the diversity of interest is not between a submitted response and example responses but between all submitted responses at a given ‘time point’ (i.e., trial number). The question is if the diversity increases or decreases as the experiment goes on and if this rate of change differs by condition. Here are the mechanics of our process. See SI Section 8.3 for more details.
@@
Appendix C Effects of Exposure and Disclosure on Creativity and Diversity
C.1. Factor 1: LLM Exposure
C.1.1. Effects on Creativity
Intuitively, the effect of exposure to ChatGPT ideas will depend on how creative ChatGPT answers are relative to human ideas. In preliminary testing, we found that the answers to the AUT generated by our prompt were scored as more creative than the ideas generated by humans (see SI 4) via the Organisciak et al. (2023) classifier. LLM generations may be increasing in creativity: while GPT-3 (an earlier model than ChatGPT-3.5) scored lower in AUT creativity than humans on the AUT (Stevenson et al., 2022), GPT-4 (a more recent model than ChatGPT-3.5) scored among the top percentile of humans on a similar verbal creativity task (Guzik et al., 2023), as measured by human judges.
C.1.2. Effects on Diversity
C.2. Factor 2: LLM Disclosure
C.2.1. Effects on Creativity
C.2.2. Effects on Diversity
It is not clear how knowing content is from AI will affect the diversity of ideas participants produce. But prior work suggests heterogeneity along two lines: the difficulty of the prompt [^8] We measure the ‘difficulty’ of a prompt by the inverse rank of the average creativity in the control condition. If participants tended to submit lower creativity ideas in the control condition for item X, we said item X was difficult., and the attitude of the participant. Prior work suggests that disclosing ideas as AI-generated would decrease diversity due to automation bias, the tendency to over-rely on AI systems (Schemmer et al., 2022; Mosier et al., 1996; Goddard et al., 2014). Increased reliance on AI ideas (when labeled as such) could lead to lower idea diversity and higher AI adoption. Conversely, some evidence suggests people display algorithmic aversion to creative products such as haikus (Hitsuwari et al., 2022) or art (Kirk et al., 2009). This aversion would yield the opposite prediction. Roemmele (2021) found that seeing AI examples only affected the participant’s writing on a key measure for difficult prompts—suggesting creative task difficulty might moderate the effect of disclosure on AI adoption. Task confidence decreases reliance on automated systems and trust in a system increases reliance on automated systems (Goddard et al., 2014). Although this literature is not usually applied to creativity, we might then suspect that self-reported creativity (i.e., roughly analogous to task confidence) and perceived AI creativity could affect the adoption of AI ideas when the source is disclosed.
Appendix D Experiment Screenshots
Appendix E Supplementary Information
SI 1. AUT Items
@@ Appendix E Supplementary Information
Table 4. AUT items by frequency of occurrence in dataset and classifier accuracy. Accuracy is defined as the correlation between human ratings of creativity and model predictions. The overall accuracy was r = 0.81. The accuracy of responses from the best-performing 5-item subset was r= 0.90. The data and model are from Organisciak et al. (2022).
@@
Table 5. Summary statistics of AUT prompt experiment. Human ideas are from Organisciak et al. (2022) and include only those ideas in response to the chosen AUT items. Note that in some cases, ChatGPT did not return the desired number of ideas, leading to a slight discrepancy between the ideas generated by the two prompts.
@@
SI 2. Pre-Treatment Questions
Pew asked about feeling towards AI (Nadeem, 2023, 2022) and we used the specific phrasing and choice ordering from (Nadeem, 2022). We randomized the first two options and kept neutral last. Our gender question was based on guidance from Spiel et al. (2019). The options were: ‘woman’, ‘man’, ‘non-binary’, ‘prefer to self-describe’, ‘prefer not to disclose’. We added a text box meant for those who preferred to self-describe. The only deviation from Spiel et al. (2019) is that we did not allow participants to select multiple options. We note that gender (as well as age and country) was optional.
SI 3. Exclusion Criteria for Analysis
Participants could have consented and answered pre-treatment questions but failed to complete any trial. We only analyze data from participants who completed at least one trial. In $(n=4)$ cases, users submitted ages that were implausible. We replaced these age values with missing for the purpose of summarizing participants but kept the responses. In $(n=2)$ cases, responses that should not have been shown were shown. We remove these two responses from the analysis. As discussed in SI SI 11, we instituted content moderation after receiving several troll responses. After the study, we manually inspected each response flagged by our system. There were 464646 ideas labeled as profane, and we determined 363636 were true positives. We remove the true positives $(n=36)$ from analysis, resulting in a final set of 341434143414 responses for analysis from an initial set of 345234523452 responses. Importantly, we conducted chi-squared tests and found that condition was unrelated to the number of flagged ideas ($\chi^{2}$(4) = 6.06, p = 0.19), number of flagged ideas minus false positives ($\chi^{2}$(4) = 2.92, p = 0.57) or total number of excluded ideas ($\chi^{2}$(4) = 3.87, p = 0.42).
SI 4. Human vs AI Ideas
We compared a sample of 1500 ideas from our modified Stevenson prompt in the prompt experiment and a random sample of 1500 ideas from the Organisciak Dataset for our 5 items. For each set, we used the model’s predicted originality scores. Originality ranges from 1-5. Overall, ChatGPT ideas had higher ($\beta=0.62$, $t(2994)=22.49$, 95% CI = [0.56,0.67][0.56,0.67][ 0.56 , 0.67 ]) originality.
Table 6. Comparing predicted originality of ChatGPT-generated ideas to ideas from a dataset of prior human responses
@@
Additionally, we wanted to minimize attrition for participants once they started. We gave participants two pieces of feedback after each trial so they would continue taking the study.
@@
SI 8. Idea Diversity
SI 8.1. Global
For 50 Monte Carlo runs with different seed values, we sampled 50 ideas for each {[condition], item} combination. For each 50-idea set, we computed various idea diversity measures. First, we calculated all pairwise SBERT distances. Next, we measured the mean and median pairwise distances. We also computed the centroid of each 50-idea set and calculated the mean distance from the centroid. After calculating these metrics, we conducted two-tailed, paired permutation tests (10,000 iterations) to test if two conditions differed on these metrics. To conduct the paired permutation test, we randomly swapped the sign of the difference between pairs of values, equivalent to randomly swapping the condition labels of rows—simulating the null hypothesis that conditions do not differ. We then counted the proportion of null distribution iterations where one would observe a larger absolute difference in means than the observed difference. We added a 1 to the numerator and denominator, which is a common, conservative adjustment (Ojala and Garriga, 2009) and stops p-values from being 0. Because the test is paired (equivalent to swapping the condition label within each ‘row’), our permutation tests are controlling for both AUT items and Monte Carlo seeds, since each row shares these attributes. We controlled for multiple pairwise comparisons by applying a Holm-Bonferroni adjustment to p-values. As a non-parametric measure of effect size, we used Cliff’s Delta. This metric ranges from -1 to +1 where 0 indicates no difference between conditions, +1 indicates that all Monte Carlo runs for the first condition are larger than those for the second, and vice versa for -1. As with evolution, to avoid the confounding effect of conditions differing in the number of seeds, we consider ideas after the sixth trial (see SI SI 8.3 for a more detailed discussion). This is because the experiment is designed to ‘shed’ all seed examples after trial six.
@@
SI 8.2. Local
Intuitively, local idea diversity is how different a response is from the examples a participant sees. There was no main effect of condition $(F(4,19.95)=0.09,p=0.98)$. We found mixed evidence that belief in AI’s relative creativity moderates local idea diversity. Regression results showed a small but significant interaction effect between the [High Exposure, Undisclosed] condition and relative AI creativity ($\beta=0.038$, $t(3244)=2.149$, 95% CI = [0.003,0.073][0.003,0.073][ 0.003 , 0.073 ], $p=0.03$). We probed this effect with estimated marginal means, predicting local idea diversity for the bottom and top decile of participants by perception of AI creativity. Top-decile participants had slightly higher local idea diversity than bottom-decile participants in the [High Exposure, Undisclosed] condition ($\Delta=2.62,d=0.34$) but although this difference was significant before multiple comparisons ($p=0.01$), it was not significant after adjusting for multiple comparisons, ($p=0.06$; see SI Table 16). Hence, we conclude there is mixed evidence for the role of belief in AI’s relative creativity as a moderator of local idea diversity.
@@
We took two additional steps to make sure our results were not driven by confounding factors. First, $Nobs$ controls for how many ideas are in the set that is being analyzed. Recall that we designed the experiment so that each (item, condition) combination was replicated exactly seven times in response chains of exactly 20 trials. However, there were some minor deviations (discussed in SI SI 11) in response chains, resulting in some (item, condition, trial number) sets having fewer items than others. Hence, we control for the number of ideas in a set. Second, we only ran this analysis on data after the sixth trial in a response chain to rule out the effect of seeds on evolution. The logic here is that the condition with the most initial seeds (6) was the control condition. The experiment is designed to ‘shed’ all seeds after trial six since by that time there would have been six experiment responses, meaning the most recent six ideas in the control condition would now all be from the experiment, and hence no seeds present in the example sets. (Note that for all local analyses, we directly control for the number of seeds present in the example set as a fixed effect.)
@@
SI 10. AI Adoption
We measured AI adoption by the maximum cosine similarity between a participant’s response and AI examples the participant saw. There was a main effect of condition $(F(3,16.59)=4.33,p=0.02)$. But because we are measuring maximum similarity, we would expect higher similarity to AI ideas in the high-exposure condition by chance (since there are more AI ideas), so we do not interpret main effects in the main manuscript and instead focus on subgroup differences.
@@
Table 20. Estimated marginal means contrasts of AI adoption, using a mixed model to compare predictions for the top 10 percentile and bottom 10 percentile of participants by belief in relative AI creativity. This metric captures how creative participants think AI is relative to humans (higher values mean more creative than humans). AI Adoption is the max cosine similarity of a participant’s response and AI examples. P-values adjusted for multiple comparisons using the Holm-Bonferroni method.